OpenStudy (anonymous):

I need help understanding how to add and subtract rational expressions. Can someone please help me?

OpenStudy (midhun.madhu1987):

Can u give me an example? i'll help u

OpenStudy (nubeer):

Probably by taking LCM

OpenStudy (anonymous):

\[\frac{ x+1 }{ x-1 }+\frac{ -2 }{ x^2-x }\]

OpenStudy (kirbykirby):

Notice that you can factor the denominator of the 2nd fraction as \(x^2-x=x(x-1)\) Now you have: \[\frac{ x+1 }{ x-1 }+\frac{ -2 }{ x(x-1) }\] If you want to add fractions, they must be on the same denominator... so you can so that by multiplying by \(\large \frac{x}{x}\) to the first fraction: \[\frac{ x+1 }{ x-1 }\frac{x}{x}+\frac{ -2 }{ x(x-1) } \\ ~ \\ =\frac{ x(x+1) }{ x(x-1) }+\frac{ -2 }{ x(x-1) } \\ ~ \\ =\frac{x(x+1)-2}{x(x-1)}\]

OpenStudy (kirbykirby):

you can do that*

OpenStudy (anonymous):

Thanks I think i got it know.

OpenStudy (anonymous):

now*